# ResFlow: Fine-tuning Residual Optical Flow for Event-based High Temporal Resolution Motion Estimation

### v1

@@ Abstract
Event cameras hold significant promise for high-temporal-resolution (HTR) motion estimation. However, estimating event-based HTR optical flow faces two key challenges: the absence of HTR ground-truth data and the intrinsic sparsity of event data. Most existing approaches rely on the flow accumulation paradigms to indirectly supervise intermediate flows, often resulting in accumulation errors and optimization difficulties. To address these challenges, we propose a residual-based paradigm for estimating HTR optical flow with event data. Our approach separates HTR flow estimation into two stages: global linear motion estimation and HTR residual flow refinement. The residual paradigm effectively mitigates the impacts of event sparsity on optimization and is compatible with any LTR algorithm. Next, to address the challenge posed by the absence of HTR ground truth, we incorporate novel learning strategies. Specifically, we initially employ a shared refiner to estimate the residual flows, enabling both LTR supervision and HTR inference. Subsequently, we introduce regional noise to simulate the residual patterns of intermediate flows, facilitating the adaptation from LTR supervision to HTR inference. Additionally, we show that the noise-based strategy supports in-domain self-supervised training. Comprehensive experimental results demonstrate that our approach achieves state-of-the-art accuracy in both LTR and HTR metrics, highlighting its effectiveness and superiority.
@@ I Introduction
Optical flow serves as a cornerstone in numerous computer vision tasks, capturing the motion of pixels over time. Recent advancements in learning-based approaches have significantly improved optical flow estimation [5, 26]. Frameworks [27, 15] incorporating all-pairs correlation computation and iterative refinement have emerged as pivotal breakthroughs, achieving remarkable performance across diverse scenarios. Despite these advancements, significant challenges persist, including the difficulty of handling high dynamic range scenes and the limitations imposed by the frame rates of conventional cameras, which restrict motion perception under complex conditions.
@@
Event cameras present a promising solution to these limitations [6]. Unlike traditional cameras, event cameras operate asynchronously, with individual pixels independently detecting changes in brightness. This design enables ultra-low latency and an exceptionally high dynamic range, making event cameras particularly effective in extreme lighting conditions and rapid motion. These advantages position event cameras as a powerful tool for motion estimation across a broader range of real-world scenarios. By leveraging the complementary strengths of event and frame-based cameras, substantial progress has been achieved in tasks such as motion deblurring [19, 38, 40], image reconstruction [20], and object tracking [29, 39]. Additionally, event-based approaches have demonstrated competitive performance compared to traditional frame-based methods [2, 3, 32, 37, 4]. However, the potential of event cameras for high-frequency applications remains underexplored. The inherently high temporal resolution of event cameras offers an unprecedented opportunity for estimating continuous motion with high precision, paving the way for transformative advancements in various downstream applications [18, 17, 41, 30, 1]. Consequently, high-temporal-resolution optical flow emerges as a promising direction in event-based vision, potentially redefining motion analysis in dynamic and challenging environments.
@@
However, the absence of real-world HTR ground truth for optical flow remains a significant bottleneck. The domain gap between synthetic datasets and real-world scenarios further limits the practical applicability of existing algorithms. Several methods have been proposed to address challenges in HTR optical flow estimation, broadly classified into two categories: (i) self-supervised methods, which use motion compensation losses to guide the estimation of intermediate motion, and (ii) cumulative methods, which progressively accumulate HTR flows to form LTR flows, indirectly supervising HTR flows with LTR ground truth. Self-supervised methods [21] heavily depend on contrast maximization loss [7, 36], leading to performance that significantly lags behind supervised approaches. In contrast, cumulative methods [35, 22] suffer from the absence of explicit constraints on intermediate flows, resulting in substantial error accumulation and ultimately limiting their effectiveness. The inherent sparsity and noise of event data become more pronounced at higher temporal resolutions, further intensifying the challenges of event-based motion estimation.
@@
To address this challenge, we propose a two-stage residual-based framework for HTR optical flow estimation and a corresponding training strategy based on LTR ground truth. Unlike cumulative methods, our framework predicts the residual flow between linear and nonlinear motion. The residual prediction requires only local robustness in correlation features, mitigating the adverse effects of sparse event data. Additionally, our training strategy resolves the discrepancies between LTR and HTR residuals, enabling effective supervision using LTR ground truth. The proposed residual-based framework addresses the inherent trade-off between temporal resolution and accuracy in cumulative methods, offering a novel and robust solution for event-based HTR optical flow estimation.
Specifically, we propose a novel residual-based paradigm that supports LTR supervision and HTR optical flow estimation. Linear motion derived from the global stage provides a robust and accurate reference for refining nonlinear motion. Residual flow prediction across varying temporal spans is unified and implemented via a shared residual refiner, enabling both LTR supervision and HTR optical flow estimation. To facilitate effective supervision of HTR flow predictions using LTR ground truth, we introduce two strategies: optical flow velocity transformation and noise-based training. In particular, we introduced regional noise that emulates residual flow patterns, facilitating the adaptation from LTR to HTR residual flows. With these learning strategies, our method utilizes ground truth with a frequency of 10 Hz for supervision, while performing inference at a frequency 15×15$\times$15 × higher (150 Hz). Extensive experiments demonstrate that our algorithm achieves state-of-the-art performance in both end-point-error and flow-warp-loss metrics. Comprehensive ablation studies further validate the effectiveness of the proposed strategies.
@@ II-B Event-based Optical Flow
Low Temporal Resolution. Event-based LTR optical flow methods primarily adapt architectures successful in frame-based optical flow. For instance, EV-FlowNet [36] voxelizes events into frame-like representations and employs the FlowNet [5] architecture for self-supervised flow estimation. As RAFT [27] gained prominence for frame-based optical flow, E-RAFT [9] adapted its design to event-based vision. Inspired by PWC-Net [26], IDNet [31] iteratively warps events and estimates optical flow, avoiding the computational overhead of constructing correlation volumes. Some methods exploit the unique characteristics of events, such as their high temporal resolution and spatial sparsity. ECDDP [34] proposed a self-supervised framework for dense prediction tasks, leveraging large-scale training on synthetic datasets. TMA [16] integrates intermediate motion information based on the correlation architecture to achieve high-quality LTR optical flow estimation. In addition to data-driven methods, model-based approaches have been explored. MultiCM [25], for instance, warps events along optical flow trajectories and formulates an energy function based on the image of warped events to handle complex scenarios. LTR methods integrate intermediate motion cues to improve the estimation accuracy of the overall trajectory, which our approach leverages as a robust reference for HTR prediction.
Figure 1: Our HTR residual flow prediction framework. HTR optical flow estimation is decoupled into two stages. We begin by splitting events, extracting features, and computing correlations to construct the temporally dense cost volumes, which serve as a shared foundation for both stages. In the global stage, all intermediate motion features are aggregated to estimate accurate LTR optical flow. In the residual stage, HTR residuals are predicted based on individual intermediate features. The shared use of cost volumes, motion encoders, and feature enhancement modules across both stages greatly reduces the model complexity. Additionally, by sharing the residual refiner across different intermediate times, the algorithm supports LTR supervision and HTR inference.
@@ III Preliminary and Event Representation
Event cameras record changes in pixel intensity asynchronously. An event $e_{i}$ is triggered when the logarithmic intensity change at a pixel exceeds the threshold $C$. Each event $e_{i}$ includes a timestamp $t$, spatial coordinates $(x_{i},y_{i})$, and a polarity $p$, where $p=+1$ indicates an increase in intensity, and $p=-1$ represents a decrease. Due to the extremely high temporal resolution of events, they are often discretized into several temporal bins for processing [36]. Specifically, events are embedded into a 3D grid representation $\mathbf{V}$ with $B$ bins as follows:
@@
In this study, to construct correlation volumes for intermediate flow, we split the event stream in the interval of $[T_{k},T_{k+1}]$ into a series of target segments $\{\mathbf{V}_{1},\mathbf{V}_{2},...,\mathbf{V}_{N}\}$ uniformly, each with an average time interval of $\Delta_{t}$. To estimate the optical flow from $T_{k}$ to $T_{k+1}$, we use the segment from time interval of $[T_{k}-\Delta_{t},T_{k}]$ as the reference $\mathbf{V}_{0}$, following previous methods [16, 10]. Eventually, we split the event into a reference $\mathbf{V}_{0}$ and multiple intermediate targets $\{\mathbf{V}_{n},n\in[1,N]\}$ to estimate the high-temporal-resolution optical flow from $T_{k}$ to $T_{k+1}$.
@@ IV-A HTR Flow Estimation via Learning HTR Residual
To mitigate the impact of event sparsity and noise on HTR prediction, our framework is structured in two stages, as shown in Fig. 1. In the global stage, the LTR optical flow is predicted using the overall event information, providing a robust reference for the intermediate flows. The residual stage focuses on estimating the nonlinear motion residuals using intermediate sparse features. We proceed to detail the two stages individually.
@@
where the $\mathbf{C}_{n}\in\mathbb{R}^{HW\times HW}$ is the cost volume between the reference $\mathbf{E}_{0}$ and the $n$-th target $\mathbf{E}_{n}$. The temporally dense cost volumes $\{\mathbf{C}_{n},n\in[1,N]\}$ are shared across both stages.
Throughout the iterations of the global stage, we perform a linear lookup within temporally dense cost volumes to retrieve the similarity cost, which is further used to derive motion features for refining the flow estimation. Given the estimated optical flow $F^{(j)}$ for the $j$-th iteration, the corresponding temporal linear lookup is formulated as follows:
$\begin{split}\mathbf{F}_{n}^{(j)}=\frac{n\Delta_{t}}{T_{k+1}-T_{k}}\cdot\mathbf{F}^{(j)},n\in\{1,2,...,N\},\\ \end{split}$ (3)
@@ IV-B Scale-consistent HTR Learning Strategy
We replace the optical flow estimate with average velocity to fill the gap caused by different time intervals. As illustrated in Fig. 4, the average velocities $\{v_{n},n\in[1,~{}N]\}$ of different intermediate flows have similar magnitudes, which simplifies the inference for intermediate flows. By predicting scale-invariant average velocities instead of optical flow displacements, we effectively address the first discrepancy between LTR and HTR residual estimation. Notably, the average velocity is ultimately transformed back into optical flow for lookup and output.
Figure 4: Training and inference pipeline illustration. We propose a strategy to support LTR supervision and HTR inference, promoting the generalization of LTR supervision to HTR predictions, which can be divided into the following three parts. Region A: Optical flow is transformed into a scale-invariant average velocity, addressing magnitude disparities between flows across varying temporal spans. Region B: The proposed regional noise is incorporated during LTR training to model HTR residual patterns. Finally, in Region C: The average velocity is converted to optical flow for lookup and output. The conversion between optical flow and velocity is still necessary during inference, but noise is no longer required. Moreover, under the noise strategy, the small LTR residual $\Delta\mathbf{F}_{N}$ enables self-supervised residual training, where the LTR-GT is replaced by the initialized LTR optical flow $\mathbf{F}_{N}$.
@@
Training is performed only at LTR flow where the GT is available, while inference can be performed at any intermediate time step. Given computational constraints, we perform inference only at the target time points. Our framework supports inference up to 15×15$\times$15 × the frequency of LTR algorithms, with results presented and analyzed in Sec. V-C.
@@ IV-C Training Objectives
Supervision is performed only where the LTR ground truth is available, which is the last flow $\mathbf{F}_{N}$. We follow the standard setup of correlation-based methods to supervise the flow. The $L_{1}$ distance between the predictions and the ground truth is taken as the loss, and the supervision is performed on each output of the iterator:
@@ V Experiment
In our experiments, the global-stage parameters were initialized using pre-trained LTR models [16] and remained frozen during the residual-stage training. Training utilized all DSEC data with available LTR-GT, and evaluation was performed on the official test set. The number of intermediate targets, $N$, was set to 5, with an LTR prediction frequency of 10 Hz and an HTR frequency of 50 Hz. During testing, we increased the ResFlow frequency to 150 Hz without retraining, results are detailed in Sec. V-C. The residual refiner iteration count was set to 4, the regional noise downsampling factor $S$ was set to 6, and the probability of adding noise was set to 0.6.
@@ V-B DSEC Dataset
Quantitative evaluation on the DSEC-Flow dataset is presented in Table I, with qualitative results shown in Fig. 6. Several conclusions can be drawn from these results. First, our proposed warping method, based on flow propagation, improves the reliability of the FWL metric. As two of the most critical evaluation metrics for event-based optical flow, EPE and FWL serve distinct purposes: EPE measures trajectory endpoint errors using LTR-GT, while FWL assesses trajectory accuracy based on IWE contrast. In previous works, these metrics often exhibited significant discrepancies, as shown in Table I. Specifically, unsupervised methods achieved favorable FWL scores but showed substantial endpoint errors, whereas supervised methods performed poorly on FWL. Previous studies [25, 21] attributed this issue to the lack of precision in warping methods. By incorporating the flow propagation method from ERAFT [9], we achieved greater consistency between EPE and FWL, leading to a more reliable evaluation. In Table I, we present FWL scores for TMA [16] using both the previous and propagation-based warping methods. The results demonstrate that the propagation-based method significantly enhances FWL scores. This improved method aligns closely with the EPE metric, further highlighting the high quality of DSEC-Flow labeling [9].
Secondly, we evaluated the performance of ResFlow and compared it to various state-of-the-art methods, as summarized in Table I. These methods are categorized into LTR and HTR approaches, with their respective supervision strategies indicated in the table. Compared to other HTR methods, our approach achieves substantial improvements in both endpoint error and trajectory accuracy. Notably, while most HTR methods utilize IWE contrast for supervision, our method relies exclusively on GT supervision and outperforms existing HTR methods on the FWL metric. For LTR methods, ResFlow maintains low endpoint errors while significantly surpassing others in FWL performance. Overall, ResFlow enhances the trajectory accuracy of linear models while preserving low endpoint errors. It is worth noting that FWL differences below 0.1 are considered significant [25]. Fig. 6 provides a qualitative comparison between the LTR model and the proposed ResFlow. The LTR model struggles to capture high-frequency nonlinear trajectories, resulting in blur artifacts. ResFlow improves motion compensation by effectively predicting HTR residuals.

### v2

@@ Abstract
Event cameras offer great potential for high-temporal-resolution (HTR) motion estimation in dynamic real-world scenarios. However, the lack of dense HTR ground truth in real-world datasets prevents fully leveraging the high temporal resolution potential of event cameras. Furthermore, the intrinsic sparsity of event data introduces additional challenges for optimization and supervision. To address these issues, we propose a residual-based paradigm that decomposes HTR optical flow into a global linear component and high-frequency residuals. The residual paradigm effectively mitigates the impacts of event sparsity on optimization and is compatible with any LTR algorithm. In addition, to bridge the supervision gap caused by the lack of HTR ground truth, we incorporate novel learning strategies. Specifically, we initially employ a shared refiner to estimate the residual flows, enabling both LTR supervision and HTR inference. Subsequently, we introduce regional noise to simulate the residual patterns of intermediate flows, facilitating the adaptation from LTR supervision to HTR inference. Additionally, we show that the noise-based strategy supports in-domain self-supervised training. Comprehensive experimental results demonstrate that our approach achieves state-of-the-art accuracy among existing HTR methods, highlighting its effectiveness and superiority.
@@ I Introduction
Optical flow serves as a cornerstone in numerous computer vision tasks, capturing the motion of pixels over time. Recent advancements in learning-based approaches have significantly improved optical flow estimation [1, 2]. Frameworks [3, 4] incorporating all-pairs correlation computation and iterative refinement have emerged as pivotal breakthroughs, achieving remarkable performance across diverse scenarios. Despite these advancements, significant challenges remain, particularly in handling high dynamic range scenes and coping with the frame-rate limitations of conventional cameras, which hinder motion perception under complex conditions.
@@
Event cameras present a promising solution to these limitations [5]. Unlike traditional cameras, event cameras operate asynchronously, with individual pixels independently detecting changes in brightness. This design enables ultra-low latency and an exceptionally high dynamic range, making event cameras particularly effective in extreme lighting conditions and rapid motion. These advantages position event cameras as a powerful tool for motion estimation across a broader range of real-world scenarios. By leveraging the complementary strengths of event and frame-based cameras, substantial progress has been achieved in tasks such as motion deblurring [6, 7, 8, 9], image reconstruction [10, 11], and object tracking [12, 13]. Additionally, event-based approaches have demonstrated competitive performance compared to traditional frame-based methods [14, 15, 16, 17, 18]. However, the potential of event cameras for high-frequency applications remains underexplored. The inherently high temporal resolution of event cameras offers an unprecedented opportunity for estimating continuous motion with high precision, paving the way for transformative advancements in various downstream applications [19, 20, 21, 22, 23, 24, 25, 26]. Consequently, high-temporal-resolution optical flow emerges as a promising direction in event-based vision, potentially redefining motion analysis in dynamic and challenging environments. Related applications have also been explored in several works [27, 28, 29, 30, 31, 32, 33, 34, 35, 36].
@@
A key bottleneck in this field is the lack of HTR ground truth in real-world event-based datasets. The domain gap between synthetic datasets and real-world scenarios further limits the practical applicability of existing algorithms. Several methods have been proposed to address challenges in HTR optical flow estimation, broadly classified into two categories: (i) self-supervised methods, which use motion compensation losses to guide the estimation of intermediate motion, and (ii) cumulative methods, which progressively accumulate HTR flows to form LTR flows, indirectly supervising HTR flows with LTR ground truth. Self-supervised methods [37] heavily depend on contrast maximization loss [38, 39], leading to performance that significantly lags behind supervised approaches. In contrast, cumulative methods [40, 41] suffer from the absence of explicit constraints on intermediate flows, resulting in substantial error accumulation and ultimately limiting their effectiveness. The sparsity and noise inherent to event data are further amplified at higher temporal resolutions, making HTR flow estimation particularly challenging.
@@
To address this challenge, we propose a two-stage residual-based paradigm for HTR optical flow estimation and a corresponding training strategy based on LTR ground truth. Unlike cumulative methods, our approach predicts the residual flow between linear and nonlinear motion. By requiring only local robustness in correlation features, residual prediction helps mitigate the adverse effects of event sparsity and noise at high temporal resolutions. Additionally, our training strategy resolves the discrepancies between LTR and HTR residuals, enabling effective supervision using LTR ground truth. The proposed residual-based framework addresses the inherent trade-off between temporal resolution and accuracy in cumulative methods, offering a novel and robust solution for event-based HTR optical flow estimation.
Specifically, we propose a novel residual-based framework that supports LTR supervision and HTR optical flow estimation. Linear motion derived from the global stage provides a robust and accurate reference for refining nonlinear motion. Residual flow prediction across varying temporal spans is unified and implemented via a shared residual refiner, enabling both LTR supervision and HTR optical flow estimation. To facilitate effective supervision of HTR flow predictions using LTR ground truth, we introduce two strategies: optical flow velocity transformation and noise-based training. In particular, we introduced regional noise that emulates residual flow patterns, facilitating the adaptation from LTR to HTR residual flows. With these learning strategies, our method utilizes ground truth with a frequency of 10 Hz for supervision, while performing inference at a frequency $15\times$ higher (150 Hz). Extensive experiments validate that our algorithm achieves state-of-the-art performance across all existing HTR methods, in terms of both end-point error and flow-warp loss. Comprehensive ablation studies further validate the effectiveness of the proposed strategies.
@@ II-B Event-based Optical Flow
Low Temporal Resolution. Event-based LTR optical flow methods primarily adapt architectures successfully in frame-based optical flow. For instance, EV-FlowNet [39] voxelizes events into frame-like representations and employs the FlowNet [1] architecture for self-supervised flow estimation. As RAFT [3] gained prominence for frame-based optical flow, E-RAFT [46] adapted its design to event-based vision. Inspired by PWC-Net [2], IDNet [47] iteratively warps events and estimates optical flow, avoiding the computational overhead of constructing correlation volumes. Some methods exploit the unique characteristics of events, such as their high temporal resolution and spatial sparsity. ECDDP [48] proposed a self-supervised framework for dense prediction tasks, leveraging large-scale training on synthetic datasets. TMA [49] integrates intermediate motion information based on the correlation architecture to achieve high-quality LTR optical flow estimation. In addition to data-driven methods, model-based approaches have been explored. MultiCM [50, 51], for instance, warps events along optical flow trajectories and formulates an energy function based on the image of warped events to handle complex scenarios. LTR methods integrate intermediate motion cues to improve the estimation accuracy of the overall trajectory, which our approach leverages as a robust reference for HTR prediction.
Figure 1: Our HTR residual flow estimation framework. HTR optical flow estimation is decoupled into two stages. We begin by splitting events, extracting features, and computing correlations to construct the temporally dense cost volumes, which serve as a shared foundation for both stages. In the global stage, all intermediate motion features are aggregated to estimate accurate LTR optical flow. In the residual stage, the HTR residual flow at each intermediate time is predicted using its corresponding motion features. The shared use of cost volumes, motion encoders, and feature enhancement modules across both stages greatly reduces the model complexity. Additionally, by sharing the residual refiner across different intermediate times, the algorithm supports LTR supervision and HTR inference.
@@ III Preliminary and Event Representation
Event cameras record changes in pixel intensity asynchronously. An event $e_{i}$ is triggered when the logarithmic intensity change at a pixel exceeds the threshold $C$. Each event $e_{i}$ includes a timestamp $t$, spatial coordinates $(x_{i},y_{i})$, and a polarity $p$, where $p=+1$ indicates an increase in intensity, and p=1p=-1 represents a decrease. Due to the extremely high temporal resolution of events, they are often discretized into several temporal bins for processing [39]. Specifically, events are embedded into a 3D grid representation $\mathbf{V}$ with $B$ bins as follows:
@@
In this study, to construct correlation volumes for intermediate flow, we split the event stream in the interval of $T_{k},T_{k+1}]$ into a series of target segments $\{\mathbf{V}_{1},\mathbf{V}_{2},...,\mathbf{V}_{N}\}$ uniformly, each with an average time interval of $\Delta_{t}$. To estimate the optical flow from $T_{k}$ to $T_{k+1}$, we use the segment from time interval of $T_{k}-\Delta_{t},T_{k}]$ as the reference $\mathbf{V}_{0}$, following previous methods [49, 55]. Eventually, we split the event into a reference $\mathbf{V}_{0}$ and multiple intermediate targets $\{\mathbf{V}_{n},n\in[1,N]\}$ to estimate the high-temporal-resolution optical flow from $T_{k}$ to $T_{k+1}$.
@@ IV-A HTR Flow Estimation via Learning HTR Residual
To mitigate the impact of event sparsity and noise on HTR prediction, our framework is structured in two stages, as shown in Fig. 1. In the global stage, the LTR optical flow is predicted using the overall event information, providing a robust reference for the intermediate flows. The residual stage focuses on estimating nonlinear residual flows at intermediate times using temporally local motion features. We proceed to detail the two stages individually.
@@
where the $\mathbf{C}_{n}\in\mathbb{R}^{HW\times HW}$ is the cost volume between the reference $\mathbf{E}_{0}$ and the $n$-th target $\mathbf{E}_{n}$, and the $D$ is the channel of feature $\mathbf{E}$. The temporally dense cost volumes $\{\mathbf{C}_{n},n\in[1,N]\}$ are shared across both stages.
Throughout the iterations of the global stage, we perform a linear lookup within temporally dense cost volumes to retrieve the similarity cost, which is further used to derive motion features for refining the flow estimation. Given the estimated optical flow $\mathbf{F}^{(j)}$ for the $j$-th iteration, the corresponding temporal linear lookup is formulated as follows:
[…]
@@ IV-B Scale-consistent HTR Learning Strategy
We replace the optical flow estimate with the average velocity to fill the gap caused by different time intervals. As illustrated in Fig. 4, the average velocities $\{v_{n},n\in[1,~N]\}$ of different intermediate flows have similar magnitudes, which simplifies the inference for intermediate flows. By predicting scale-invariant average velocities instead of optical flow displacements, we effectively address the first discrepancy between LTR and HTR residual estimation. Notably, the average velocity is ultimately transformed back into optical flow for lookup and output.
Figure 4: Training and inference pipeline illustration. We propose a strategy to support LTR supervision and HTR inference, promoting the generalization of LTR supervision to HTR predictions, which can be divided into the following three parts. Region A: Optical flow is transformed into a scale-invariant average velocity, addressing magnitude disparities between flows across varying temporal spans. Region B: The proposed regional noise is incorporated during LTR training to model HTR residual patterns. Here, -1 indicates that the estimated noise is subtracted from the noise-perturbed flow. Finally, in Region C: The average velocity is converted to optical flow for lookup and output. The conversion between optical flow and velocity is still necessary during inference, but noise is no longer required. Moreover, under the noise strategy, the small LTR residual $\Delta\mathbf{F}_{N}$ enables self-supervised residual training, where the LTR-GT is replaced by the initialized LTR optical flow $\mathbf{F}_{N}$.
@@
Training is performed only at LTR flow where the GT is available, while inference can be performed at any intermediate time step. Given computational constraints, we perform inference only at the target time points. Our framework supports inference up to $15\times$ the frequency of LTR algorithms, with results presented and analyzed in Sec. V-C.
@@ IV-C Training Objectives
Supervision is performed only where the LTR ground truth is available, which is the last flow $\mathbf{F}_{N}$. We follow the standard setup of correlation-based methods and adopt the $L_{1}$ distance between the predictions and the ground truth as the loss function, due to its robustness to outliers and stable gradient properties. The supervision is applied to each output of the iterator.
@@ V Experiment
In our experiments, the global-stage parameters were initialized using pre-trained LTR models [49] and remained frozen during the residual-stage training. Training utilized all DSEC data with available LTR-GT, and evaluation was performed on the official test set. The number of intermediate targets, $N$, was set to 5, with an LTR prediction frequency of 10 Hz and an HTR frequency of 50 Hz. During testing, we increased the ResFlow frequency to 150 Hz without retraining, and results are detailed in Sec. V-C. The residual refiner iteration count was set to 4, the regional noise downsampling factor $S$ was set to 6, and the probability of adding noise was set to 0.6.
@@ V-B DSEC Dataset
Table I presents the quantitative evaluation on the DSEC-Flow dataset, while Fig. 6 shows the qualitative results. To ensure a comprehensive assessment, we report results for both LTR and HTR methods, along with the corresponding LTR (EPE and %Out) and HTR (FWL) metrics. Several conclusions can be drawn from these results. First, our proposed warping method, based on flow propagation, improves the reliability of the FWL metric. As two of the most critical evaluation metrics for event-based optical flow, EPE and FWL serve distinct purposes: EPE measures trajectory endpoint errors using LTR-GT, while FWL assesses trajectory accuracy based on IWE contrast. In previous works, these metrics often exhibited significant discrepancies, as shown in Table I. Specifically, unsupervised methods achieved favorable FWL scores but showed substantial endpoint errors, whereas supervised methods performed poorly on FWL. Previous studies [50, 37] attributed this issue to the lack of precision in warping methods. By incorporating the flow propagation method from ERAFT [46], we achieved greater consistency between EPE and FWL, leading to a more reliable evaluation. In Table I, we present FWL scores for TMA [49] using both the previous and propagation-based warping methods. The results demonstrate that the propagation-based method significantly enhances FWL scores. This improved method aligns closely with the EPE metric, further highlighting the high quality of DSEC-Flow labeling [46].
Secondly, we evaluated the performance of ResFlow and compared it to various state-of-the-art methods, as summarized in Table I. These methods are categorized into LTR and HTR approaches, with their respective supervision strategies indicated in the table. Compared to other HTR methods, our approach achieves substantial improvements in both endpoint error and trajectory accuracy. Notably, while most HTR methods utilize IWE contrast for supervision, our method relies exclusively on GT supervision and outperforms existing HTR methods on the FWL metric. For LTR methods, ResFlow maintains low endpoint errors while significantly surpassing others in FWL performance. Overall, ResFlow enhances the trajectory accuracy of linear models while preserving low endpoint errors. It is worth noting that FWL differences below 0.1 are considered significant [50]. Fig. 6 provides a qualitative comparison between the LTR baseline and the proposed ResFlow. The LTR model struggles to capture high-frequency nonlinear trajectories, particularly around object edges and fast motions, leading to noticeable blur and distortion artifacts. In contrast, ResFlow enhances motion compensation by accurately predicting HTR residuals, resulting in sharper and more temporally consistent IWE.